OpenStudy (anonymous):

How many real number solutions does the following equation have? –8x^2 – 8x – 2 = 0

OpenStudy (inkyvoyd):

0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Do you understand how to get that answer?

OpenStudy (anonymous):

There are two real zeros. They are multiples of x=-1/2 So x=-1/2 is a solution.\[(x+1/2)^{2}\]Check \[-8(-1/2)^{2}+-8(-1/2)-2=0\]\[-8(1/4)+4-2=0\]\[-2+4-2=0\]\[0=0\]